Pentane is burnt with 96000 cm3 of oxygen gas at room temperature and pressure.
Calculate the mass of zinc needed to react with 50 cm3 of 0.8 M of Hydrochloric acid.
1
Expert's answer
2021-10-12T07:48:18-0400

96000 cm3 of oxygen is "\frac{96000}{1000\times22.4}=4.29" mol of oxygen

The reaction between pentane and oxygen is:

C5H12 + 8O2 = 5CO2 + 6H2O


50 cm3 of 0.8 M HCl contains "\frac{50\times0.8}{1000} = 0.04 \ mol" of HCl

The equotion between Zn and HCl:

Zn + 2HCl = ZnCl2 + H2

The mass of zink needed is "\frac{0.04\times65.38}{2} = 1.3076" g